The following information is available for Lock-Down Company, which produces special-order security
products and uses a job order cost accounting system.
April 30 May 31
Inventories
Raw materials . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . $40,000 $ 50,000
Goods in process . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9,600 19,500
Finished goods . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60,000 33,200
Activities and information for May
Raw materials purchases (paid with cash) . . . . . . . . . . . . . . . . . . . . . . 189,000
Factory payroll (paid with cash) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 400,000
Factory overhead
Indirect materials . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12,000
Indirect labor . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75,000
Other overhead costs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 100,500
Sales (received in cash) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1,200,000
Predetermined overhead rate based on direct labor cost . . . . . . . . . . 65%
Compute the following amounts for the month of May.
1. Cost of direct materials used. 4. Cost of goods sold.*
2. Cost of direct labor used. 5. Gross profit.
3. Cost of goods manufactured. 6. Overapplied or underapplied overhead.
*Do not consider any underapplied or overapplied overhead.
Check (3) $693,350
Answer #1

1) Cost of direct material used = 40000+189000-50000-12000 = 167000

2) Cost of direct labor used = 400000-75000 = 325000

3) Cost of goods manufactured = 167000+1.65*325000+9600-19500 = 693350

4) Cost of goods sold = 60000+693350-33200 = 720150

5) Gross profit = 1200000-720150 = 479850

6) Applied overhead = 325000*65% = 211250

Actual overhead = 187500

Over applied overhead = 211250-187500 = 23750